If a baseball player's batting average is 0.340 ( i.e., the probability of getting a hit each time at bat is 0.340), find the probability that the player will have a bad season and get at most 60 hits in 200 times at bat?


1
Expert's answer
2021-05-07T14:32:24-0400

We have given that,

"n = 200"

"p = 0.34"

"x=" number of hits

"n=" total times at bat

We know,


"\mu = np = 200 \times 0.34 = 68"


"\sigma = \sqrt{npq} = \sqrt{200 \times 0.34 \times 0.66} = 6.6993"


Applying continuity correction,


"z = \dfrac{x-\mu}{\sigma}"


"z = \dfrac{60-68}{6.6993} = -1.12"


For the normal standard distribution,


"P(z < - 1.12) = 13.14% = 13.1" %